a clerk counts $1.00 and $10.00 bills at the end of the day he finds he has a total of 49 bills having a combined monetary value of $139 . find the number of bills of each denomination
the clerk has --------ones and ------------tens?

Answer by amoresroy(361)

a clerk counts $1.00 and $10.00 bills at the end of the day he finds he has a total of 49 bills having a combined monetary value of $139 . find the number of bills of each denomination
Let x = the number of $1.00 bills
49-x = the number of $10.00 bills
1.00x + 10.00(49-x) = 139
x + 490 -10x = 139
Combine like terms
-9x = 139-490
x = -351/-9
x = 39
49-x = 10
The clerk has 39 $1 bills and 10 $10 bills